10. Find the intensity of a magnetic field at a point 2 sm away from an infinitely long conductor which a current of 5 A flows through.


1
Expert's answer
2022-01-31T10:25:22-0500

Magnetic field

"B=\frac{\mu_0 I}{2\pi r}"

Put value

I=5A

r=2cm =0.02m

"B=\frac{4\times3.14\times\times10^{-7}\times5}{2\pi \times0.02}=5\times10^{-5} T"